OpenStudy (anonymous):

Find the remainder when f(x) = 2x^3 – 12x^2 + 11x + 2 is divided by x – 5. is the answer 7? I'm not sure?

OpenStudy (lgbasallote):

according to wolframalpha it's not

OpenStudy (lgbasallote):

you may have done something wrong in your division

OpenStudy (anonymous):

okay. I dnt know where I went wrong then...

OpenStudy (anonymous):

did u use synthetic division??? i would

OpenStudy (lgbasallote):

or i inputted something wrong in wolfram

OpenStudy (lgbasallote):

im gonna do this by hand now..

OpenStudy (anonymous):

No I didn't use that method.

OpenStudy (anonymous):

–3 3 –7

OpenStudy (anonymous):

its all ways eaiser

OpenStudy (anonymous):

these are the rest of the answers.

OpenStudy (anonymous):

what did you get on wolfalpha?

OpenStudy (lgbasallote):

wait..this software is so choosy lol

OpenStudy (anonymous):

i got -7

OpenStudy (lgbasallote):

lol i hate this software

OpenStudy (lgbasallote):

yeah i think it's -7

OpenStudy (anonymous):

okay lol. and thank you katie.

OpenStudy (anonymous):

ur welcome!!

OpenStudy (lgbasallote):

moral: dont trust a software haha

OpenStudy (anonymous):

and yah it really is choosy so that's why I dnt use it.

OpenStudy (anonymous):

ikr :)

OpenStudy (anonymous):

yeah i prefer free hand lol

OpenStudy (anonymous):

I'll try that one next time lol well thanks guys.

OpenStudy (anonymous):

\[f(x) = 2x^3 – 12x^2 + 11x + 2 \]\[f(5) = 2\times 5^3 – 12x\times 5^2 + 11\times5 + 2 \] is your remainder

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

\[f(5) = 2\times 5^3 – 12\times 5^2 + 11\times5 + 2\] typo there

OpenStudy (lgbasallote):

@satellite73 remind me again how to use remainder theorem?

OpenStudy (anonymous):

very sure

OpenStudy (anonymous):

if you divide a polynomial \(p(x)\) by \(x-k\) your remainder is \(p(k)\)

OpenStudy (anonymous):

that's not any of my choices tho, so Idk.

OpenStudy (anonymous):

did you compute?

OpenStudy (anonymous):

no, I re did it by hand.

OpenStudy (anonymous):

and I got -7 like they said.

OpenStudy (anonymous):

i get 7

OpenStudy (lgbasallote):

yup i got 7 too using remainder theorem

OpenStudy (lgbasallote):

how'd you make wolframalpha work o.O

OpenStudy (anonymous):

ahhhh sorry

OpenStudy (anonymous):

i used the rong sign

OpenStudy (lgbasallote):

oh lol i made a syntax error in wolfram a while ago no wonder

OpenStudy (anonymous):

\[f(x)=2 x^3-12 x^2+11 x+2 = (1-2 x+2 x^2)×(x-5)+7\] and so it is pretty clear that \(f(5)=7\) because the second factor will be 0 leaving 7 as the answer